### v1

# VersaTune: Harnessing Vertical Domain Insights for Multi-Ability LLM Supervised Fine-Tuning

Abstract
Large Language Models (LLMs) exhibit remarkable capabilities in handling multiple tasks across domains due to their emergent properties. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-ability performances during fine-tuning. We categorize knowledge into distinct domains including law, medicine, finance, science, code. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the composition of training data that aligns with the model’s existing knowledge distribution. During the fine-tuning process, weights of different domains are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results demonstrate that VersaTune achieves significant improvements in multi-domain performance, with a enhancement in comprehensive multi-domain tasks. Additionally, in scenarios where specific domain optimization is required, VersaTune reduces the degradation of performance in other domains by , without compromising the target domain’s training efficacy.
1 Introduction
Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) (Achiam et al., 2023; Dwivedi et al., 2021; Lewkowycz et al., 2022), particularly for Natural Language Processing tasks (Brown et al., 2020; Devlin, 2018; Radford et al., 2019), reshaping AI research and applications in domains such as law (Cui et al., 2023), medical (Singhal et al., 2023; Thirunavukarasu et al., 2023), finance (Li et al., 2023; Wu et al., 2023), science (Beltagy et al., 2019; Taylor et al., 2022) and code (Liu et al., 2024; Roziere et al., 2023). In order to further incentivize capabilities in these areas, LLMs typically undergo supervised fine-tuning (SFT) stages on specific datasets.
As demonstrated by the robust performances of proprietary LLMs such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on SFT tends to merely concentrate on a single ability of LLMs (Dong et al., 2023; Xu et al., 2024). We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during the SFT phase:
How to design a data composition strategy during LLMs’ SFT stages that could achieve overall multitasking capabilities?
Through extensive experimental analysis, we identified that the challenges associated with data composition strategies stem from the following three key aspects:
C1: Catastrophic Forgetting. Given the fundamental differences between tasks across various domains, for multi-domain SFT, the sequential training strategy of specific datasets in multiple phases can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting (Kaushik et al., 2021; McCloskey & Cohen, 1989). This problem hinders the versatile fine-tuning performance of a model across multiple abilities (De Lange et al., 2021; Dong et al., 2023; Yuan et al., 2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during the SFT stage, enhancing the overall performance and adaptability.
C2: Low Efficiency. Existing research on data composition during the SFT phase for LLMs is still in its initial stages, with most strategies based on heuristic rules (Albalak et al., 2024; Dubey et al., 2024; Wang et al., 2023). Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience (Azerbayev et al., 2023; Roziere et al., 2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of domains during the current phase based on real-time feedback from the LLMs, which hampers the ability to minimize the loss of performances in other domains as the LLMs undergo specialized training.
To address these challenges, we introduce VersaTune, a novel data composition framework designed for enhancing models’ overall performances across different vertical domains during fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 2.2), followed by data composition based on the existing distribution for multi-ability enhancement (Section 3.2) as well as flexible domain expansion (Section 3.3). VersaTune exhibits the following properties:
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ fine-tuning stage, providing an efficient strategy for enhancing versatile capabilities (C2).
Flexible. VersaTune can be flexibly adapted to scenarios that expand performance on specific domain tasks while minimizing the degradation of the model’s capabilities in non-target domains (C1, C3).
Robust. The data composition strategy demonstrates significant improvements after fine-tuning in open-source models with parameter sizes ranging from 7B-14B, adding to the effectiveness of VersaTune .
We evaluate our VersaTune on 12 downstream tasks across vertical domains, including law, medicine, finance, science, and code, as well as comprehensive datasets, each of which features two benchmarks. The results show that VersaTune can achieve an improvement of in versatile fine-tuning across multiple domains. Furthermore, when focusing on specific-domain fine-tuning, VersaTune maintains the training effectiveness in the target domain while reducing the performance degradation in other non-target domains by .
2 VersaTune: Detecting Knowledge Distribution
In this section, we present a domain mixing strategy for fine-tuning a LLM that possesses a comprehensive multi-task capability (Section 2.1). This approach is designed to align with the inherent domain knowledge distribution of the base model. Subsequently, we describe the method for detecting domain knowledge proportion of the base model, which is crucial for informing the fine-tuning process (Section 2.2).
2.1 Knowledge Consistency Training
Previous research on data mixing ratios during the SFT phase for LLMs has predominantly focused on enhancing capabilities within a specific domain, often utilizing only data from that domain or employing heuristic, experience-based data proportions. We argue that such data mixing strategies can significantly impair the LLM’s abilities in other domains. In the fine-tuning stage, maintaining a robust overall capability across various domains is crucial.
What data mixing strategy effectively boosts the versatile performance of LLMs across different domains during the SFT phase? We propose the following statement:
Statement 1 An LLM fine-tuned with domain-specific data proportions that align with its pre-trained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
where denotes the set of all possible data points.
The statement is based on the fact that the model has developed an understanding of language features and knowledge specific to various domains during pretraining. Maintaining the same distribution of knowledge during fine-tuning allows the model to build on this pre-existing knowledge, enhancing learning efficiency and robustness.
2.2 Domain Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods (Gekhman et al., 2024; Zhao et al., 2023b) and training data inference strategies for LLMs (Ding et al., 2022; Hayase et al., 2024), we propose a structured approach to efficiently detect domain knowledge based on statistics.
Assuming the data corpus encompasses distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary LLM to infer the probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for the domain knowledge distribution.
Input: Base model to be fine-tuned , Proprietary model , Hyperparameters: sample number , maximum iterations
Parameter: Data samples generated from the base model
Output: Domain distribution
Define : domain probability distribution of data sample
3 VersaTune: Fine-Tuning Multi-Ability LLMs Efficiently
Having detected the distribution of domain knowledge within the base model, we will now utilize these findings to guide our multi-ability SFT process. The approaches are designed to enhance the overall performance of the fine-tuned model across a spectrum of multi-domain tasks (Section 3.2), as well as to facilitate the flexible expansion of capabilities in specific domains (Section 3.3).
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a domain and its corresponding proportion, such that training a model on could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performance of other domains. The specialized capabilities of LLMs are ultimately measured using downstream tasks related to different domains (e.g., FinBen (Xie et al., 2024a) for financial performances).
3.1 Preliminary: Learnable Potential and Forgetting Degree of Domain Knowledge
Before formally introducing the effective multi-task fine-tuning and flexible domain expansion data composing strategies, we will first provide an overview of the evaluation metrics used for the following algorithms in this subsection.
Mastery Ceiling. We first fine-tuned the small reference model for epochs on each individual domain separately, and identified the epoch with the lowest average loss during this process as the lower bound on the minimum loss attainable by the target model for the given domain. This value represents the highest level of domain knowledge mastery that the target model can achieve in the context of the current specific domain under given conditions.
Learnable Potential. We can observe whether a domain could be effectively learned by the model through comparing the difference between the loss of the target model and the minimum loss that the reference model can achieve. Based on these principles, we propose Equation 2 to score the learnable potential of domain .
where denotes the loss associated with the target model for the -th domain, while signifies the corresponding loss for the reference model within the same domain. To mitigate the impact of inherent loss variations across different domains for the model, we have introduced a normalization term into the formula.
Forgetting Degree. When focusing on expanding a model to a specific domain, our objective is to minimize the loss of the model’s knowledge regarding other domains. Here we segment the fine-tuning stage into distinct checkpoints. We quantify the degree of knowledge loss, or the forgetting of the current domain, by measuring the difference in loss between the -th and -th training steps. This difference reflects the model’s mastery loss for the tasks associated with the current domain. Based on this principle, we introduce Equation 3 to assess the model’s forgetting degree for domain at the -th training step.
where represents the loss at the -th training step associated with the target model for the -th domain, while denotes the loss at the preceding -th iteration for the same domain. We also incorporated a normalization factor into the equation to counteract the effects of inherent loss disparities among domains.
3.2 Effective Multi-Ability Fostering
Input: Base model to be fine-tuned , Domain reference loss , Hyperparameters: magnitude of adjustment , number of training steps
Parameter: Data proportion of the SFT dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
To cultivate the multi-tasking capabilities of a LLM during the fine-tuning phase, we have aligned the initial domain distribution in the SFT stage with the knowledge detection results of the base model as stated in Equation 1. Furthermore, we dynamically make minor adjustments in the composition ratios of various domains based on the model’s real-time feedback at different SFT stages.
As detailed in Algorithm 2, in the pursuit of balanced domain expertise enhancement, we first blended the domain proportions in accordance with the base model’s intrinsic domain knowledge distribution detected by Algorithm 1 (Line 1). Then at each training step , we assigned a learnable potential score to each domain based on the methodology outlined in Equation 2. These scores were then utilized to fine-tune the representation of each domain within the composite SFT dataset, ensuring a balanced development of competencies across all domains throughout the fine-tuning process (Line 3-7). At the same time, the parameters of model are updated based on the gradients computed through backpropagation (Line 8). This adaptive approach is imperative to harmonize the progression of capabilities in different domains and to optimize the model’s performance on a multifaceted array of tasks.
3.3 Flexible Domain Expansion
Input: Base model to be fine-tuned , Domains that require enhanced cultivation , Domain reference loss , Hyperparameters: number of training steps , magnitude of adjustment , extent of domain proportion adjustment , variation threshold
Parameter: Data proportion of the SFT dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
Define : forgetting degree of the current domain
When conducting fine-tuning on a pre-trained model, there are instances where we aim to particularly enhance the model’s performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
Statement 2
When fine-tuning a LLM for a specific capability, increasing the volume of data from a particular domain while adjusting other domains according to the knowledge distribution of the base model, facilitates a flexible strategy for domain expansion. Formally, the relationship can be represented as:
where is the data distribution in the given domain , and is the adjustment factor.
Here is determined based on the knowledge distribution of the pre-trained domain. In particular, when increases, the other domains shrink proportionally as a whole, which can be expressed as:
where is the increased adjustment factor, and is the original ratio of other domain knowledge relative to .
As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pre-trained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other domains: if the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other domains according to Equation 5 (Line 8-9); otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
4 Experiments
In this section, we describe details of our experimental setup (Section 4.1), the baseline methods we use for comparison (Section 4.2), and experimental results (Section 4.3).
| Model | Method | Law | Medical | Finance | Science | Code | General | Avg. |
| LLaMA-2-7B | Uniform Distribution | 23.22 | 25.51 | 18.11 | 26.10 | 4.56 | 45.78 | 23.88 |
| Inverse Distribution | 20.09↓ | 23.95↓ | 17.75↓ | 24.33↓ | 2.95↓ | 45.08↓ | 22.36↓ | |
| VersaTune | 29.75↑ | 37.90↑ | 32.66↑ | 53.41↑ | 15.65↑ | 48.22↑ | 36.27↑ | |
| LLaMA-3-8B | Uniform Distribution | 32.34 | 20.65 | 19.90 | 28.34 | 25.90 | 48.57 | 29.28 |
| Inverse Distribution | 27.66↓ | 17.14↓ | 18.51↓ | 26.36↓ | 22.01↓ | 47.93↓ | 26.60↓ | |
| VersaTune | 43.77↑ | 43.97↑ | 41.19↑ | 58.64↑ | 32.28↑ | 51.20↑ | 45.18↑ | |
| Qwen-1.5-7B | Uniform Distribution | 31.91 | 16.71 | 34.02 | 29.15 | 29.36 | 54.50 | 32.61 |
| Inverse Distribution | 29.97↓ | 13.03↓ | 33.12↓ | 27.53↓ | 26.25↓ | 54.58↓ | 30.75↓ | |
| VersaTune | 38.00↑ | 31.36↑ | 56.08↑ | 54.44↑ | 36.15↑ | 57.24↑ | 45.55↑ | |
| Qwen-2-7B | Uniform Distribution | 35.52 | 26.40 | 37.06 | 44.02 | 45.54 | 60.46 | 41.50 |
| Inverse Distribution | 30.91↓ | 20.11↓ | 33.92↓ | 38.41↓ | 42.14↓ | 59.63↓ | 37.52↓ | |
| VersaTune | 43.05↑ | 43.36↑ | 64.67↑ | 61.00↑ | 52.90↑ | 60.85↑ | 54.31↑ | |
| Qwen-2.5-7B | Uniform Distribution | 35.80 | 25.51 | 45.62 | 42.86 | 51.25 | 59.56 | 43.43 |
| Inverse Distribution | 31.67↓ | 21.76↓ | 43.16↓ | 39.73↓ | 48.80↓ | 59.30↓ | 40.74↓ | |
| VersaTune | 44.20↑ | 43.19↑ | 65.79↑ | 71.15↑ | 60.97↑ | 61.82↑ | 57.85↑ | |
| LLaMA-2-13B | Uniform Distribution | 41.26 | 34.76 | 35.20 | 48.25 | 37.55 | 55.20 | 42.04 |
| Inverse Distribution | 35.40↓ | 27.53↓ | 31.72↓ | 43.82↓ | 32.31↓ | 54.89↓ | 37.61↓ | |
| VersaTune | 48.01↑ | 46.73↑ | 47.30↑ | 63.36↑ | 46.15↑ | 57.14↑ | 51.45↑ | |
| Qwen-2.5-14B | Uniform Distribution | 45.11 | 43.28 | 64.56 | 56.60 | 70.45 | 65.4 | 57.57 |
| Inverse Distribution | 40.72↓ | 39.16↓ | 62.71↓ | 47.95↓ | 65.93↓ | 65.15↓ | 53.60↓ | |
| VersaTune | 53.59↑ | 47.70↑ | 72.26↑ | 72.88↑ | 82.06↑ | 65.75↑ | 65.71↑ |
4.1 Experimental Setup
We have collected SFT datasets spanning 6 vertical domains, including Lawyer-Instruct111https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA (Jin et al., 2020), Finance Alpaca222https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations333https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca444https://github.com/sahil280114/codealpaca and Alpaca (Taori et al., 2023), denoted as , to represent law, medical, finance, science, code and general capabilities SFT datasets respectively. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other five specific domains, keeping only the general domain instances unrelated to them.
We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series as our pretrained language models . During the fine-tuning procedure, we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800.
4.2 Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
4.3 Results
We conduct evaluations to validate the efficiency of VersaTune across different models in scenarios that encompass both effective multi-ability fostering and flexible domain expansion, the main results of which are presented in Table 1 and Figure 2. We summarize the observations below.
VersaTune is efficient across different models in both scenarios. Table 1 shows that VersaTune consistently outperforms other methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the fine-tuning phase. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Figure 2. Detailed results and analysis can be found in Appendix D.
Knowledge consistency training boosts performance. In Table 1, we present the experimental results of data composition strategies that allocate domain data inversely proportional to the detected knowledge distribution. As expected, the inverse distribution strategy yielded lower performance compared to the simplest approach of uniform distribution across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training.
5 Ablations and Analysis
Previously in Section 4, we demonstrated the effectiveness of VersaTune in enhancing multiple abilities and enabling flexible domain expansion during the model’s fine-tuning phase. In this section, we perform an in-depth analysis of VersaTune, where we ablate the components of dynamic adaptation in Algorithm 2, the criteria for determining the upper limit of domain expansion in Algorithm 3, and the impact of reference model scale employed for gauging the learnable potential and forgetting degree of the target model.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights derived from the knowledge distribution obtained from probing the target model prior to fine-tuning. Table 4 and Figure 3 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights during training by monitoring the learnable potential of each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes throughout training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial.
Establishing proportion thresholds for specific domains counts during domain expansion. We consider conducting a comparative analysis between the outcomes of VersaTune and those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3. Figure 4 shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Concurrently, it ensures gains in the capacity for the current domain of interest. We speculate that it is because by the later stages of fine-tuning, models’ proficiency in the target domain approaches saturation. Further increasing the proportion of the current domain provides diminishing returns and can lead to a significant loss of performance in other domains.
6 Related Work
Supervised Fine-Tuning of LLMs. The Supervised Fine-Tuning (SFT) stage of a Large Language Model (LLM) involves further training to refine the model’s task-solving capabilities and ensure greater alignment with human instructions (Zhao et al., 2023a). While recent research has delved into exploring fine-tuning methods for multi-task enhancement (Dong et al., 2023; Sanh et al., 2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage.
Data Reweighting for LLM Training. Data reweighting maintains full access to the entire dataset while adjusts the relative importance of each instance for various target tasks, which is essential for both pretraining and fine-tuning stages of LLMs (Wang et al., 2023). During the pretraining stage, DoReMi (Xie et al., 2024b) and DoGE (Fan et al., 2023) employ lightweight proxy models to estimate weights for different data sources, which are subsequently applied to the formal training of LLMs. Furthermore, Sheared LLaMA (Xia et al., 2023) implements an online variant of DoReMi, where the loss reference provided by the proxy model is replaced with the state of the pre-trained model from the previous training step. Additionally, ODM (Albalak et al., 2023) leverages the differences in loss or perplexity across domains within a multi-armed bandit framework to dynamically adjust the data distribution along training steps. As for the SFT phase, Dong et al. (2023) focus on enhancing the model’s math reasoning, coding, and general human-aligning abilities through a dual-stage mixed fine-tuning strategy. However, the mixing ratios for different domains rely heavily on enumeration and empirical methods, and the covered domains are not holistic. This paper provides a comprehensive overview of the model’s capabilities across various domains during the SFT stage and proposes appropriate multi-ability fine-tuning methods.
Knowledge Detection in LLMs. Investigating the knowledge contained in current LLMs is essential for guiding their subsequent training (Chang et al., 2024). The knowledge encompasses multiple dimensions, such as different domain sources and task attributes. Existing work in LLM knowledge detection primarily focuses on prompting and calibration. Directly prompting the model to generate sequences and extracting confidence scores from the model (Gekhman et al., 2024; Kadavath et al., 2022; Kuhn et al., 2023; Manakul et al., 2023) is a common strategy. However, such approaches are highly dependent on prompt design and task selection, which can introduce bias into the assessment. Other studies have attempted to infer the training data mixtures used in previous training stages of LLMs (Antoniades et al., 2024; Hayase et al., 2024; Hu et al., 2022; Ye et al., 2022). The essence of these studies is to evaluate the current knowledge state of the models and provide targeted strategies for data organization in subsequent training phases.
7 Conclusion
The versatile capability of Large Language Models (LLMs) across multiple domain tasks is of paramount importance. Existing fine-tuning methods for LLMs often focus on single domain tasks, leading to significant degradation in performance across other domains. In this paper, we introduce VersaTune, a data composition framework designed to enhance the multi-domain capabilities of models during the fine-tuning phase, based on the domain knowledge distribution of the target model. Experimental results from different models and downstream tasks across various domains demonstrate that VersaTune achieves excellent training outcomes in both scenarios of overall multi-domain enhancement and flexible expansion of specific domains.
Appendix A Details of Knowledge Distribution Detection
During the domain knowledge detection phase for our target models, we employed Qwen2.5-72B-Instruct555https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as the proprietary model . For each target model slated for fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several vertical domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 5.
Appendix B Training
B.1 Training Datasets
For training, we selected representative datasets for each domain, which exhibit significant differences in format, sentence length, and domain-specific content. These differences reflect the heterogeneity of training data across various domains during the fine-tuning stage. Further details about these datasets can be found in Table 2. Specifically, for the Alpaca dataset, which we utilize for representing the general domain, we have excluded data related to law, medicine, finance, science, and code domains to ensure the precision and authenticity of the actual domain weight.
| Dataset | # Instance | Source | # Rounds | Full |
| Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset666https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
| MedQA | 10178 | Professional Medical Board Exams | 1 | Training Portion |
| Finance Alpaca | 68912 | Alpaca, FiQA, 1.3k Pairs Generated using GPT3.5 | 1 | |
| Sonnet3.5 Science Conversations | 8835 | Scientific Conversations with Sonnet3.5 | 11.1 | |
| Code Alpaca | 20022 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | |
| Alpaca | 49,087 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | Excluding Samples of Other Domains |
B.2 Training Details
All experiments were conducted based on full fine-tuning. We employed a learning rate scheduler that incorporates linear warm-up and cosine decay, peaking at a learning rate of 2e-5. A batch size of 128 was utilized, and the training spanned 4 epochs across all selected datasets. For scenarios aimed at fostering multi-ability, we trained and assessed models including LLaMA-2-7B, LLaMA-3-8B, LLaMA-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. In the context of domain expansion, the training and evaluations were performed using the LLaMA-3-8B and LLaMA-2-13B models. The total number of samples per epoch was set to , with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for the Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Appendix C Evaluation Details
We evaluate the performance of the models on the downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 3. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset (Zhang et al., 2024) for testing.
| Domain | Benchmark | # Instance | Language | N-Shot |
| Law | LegalBench (Guha et al., 2024) | 90,394 (164 sub-tasks) | English | 1 |
| LawBench (Fei et al., 2023) | 10,000 (20 sub-tasks) | Chinese | 1 | |
| Medicine | MedQA (Jin et al., 2020) | 1,273 | English | 1 |
| MedMCQA (Pal et al., 2022) | 4,183 | English | 1 | |
| Finance | FinEval (Zhang et al., 2023) | 4,661 (34 sub-tasks) | Chinese | 1 |
| FinanceIQ (Zhang & Yang, 2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |
| Science | SciEval (Sun et al., 2024) | 15,901 | English | 1 |
| MMLU-Sci (Hendrycks et al., 2020) | 2,999 (14 sub-tasks) | English | 0 | |
| Code | HumanEval (Chen et al., 2021) | 164 | English | 0 |
| MBPP (Austin et al., 2021) | 974 | English | 0 | |
| Other (General) | AGIEval (Zhong et al., 2023) | 8,062 (20 sub-tasks) | English, Chinese | 0 |
| HellaSwag (Zellers et al., 2019) | 10,003 | English | 0 |
Appendix D More Experiment Results
D.1 Performance on Domain-Specific Tasks in Multi-Ability Fostering Scenario
We present herein the outcomes of the ablation study along with the raw scores from domain-specific benchmarks.
| Model | Method | Law | Medical | Finance | Science | Code | General | Avg. |
| LLaMA-2-7B | VersaTune Constant | 28.51↓ | 33.45↓ | 21.09↓ | 47.89↓ | 13.47↓ | 47.10↓ | 31.92↓ |
| VersaTune | 29.75 | 37.90 | 32.66 | 53.41 | 15.65 | 48.22 | 36.27 | |
| LLaMA-3-8B | VersaTune Constant | 42.80↓ | 33.74↓ | 31.60↓ | 55.41↓ | 31.55↓ | 49.48↓ | 40.76↓ |
| VersaTune | 43.77 | 43.97 | 41.19 | 58.64 | 32.28 | 51.20 | 45.18 | |
| Qwen-1.5-7B | VersaTune Constant | 36.75↓ | 27.94↓ | 46.29↓ | 47.86↓ | 34.04↓ | 56.59↓ | 41.58↓ |
| VersaTune | 38.00 | 31.36 | 56.08 | 54.44 | 36.15 | 57.24 | 45.55 | |
| Qwen-2-7B | VersaTune Constant | 39.29↓ | 40.01↓ | 50.73↓ | 56.61↓ | 51.29↓ | 60.57↓ | 49.75↓ |
| VersaTune | 43.05 | 43.36 | 64.67 | 61.00 | 52.90 | 60.85 | 54.31 | |
| Qwen-2.5-7B | VersaTune Constant | 41.99↓ | 39.96↓ | 55.71↓ | 64.00↓ | 55.63↓ | 60.82↓ | 53.02↓ |
| VersaTune | 44.20 | 43.19 | 65.79 | 71.15 | 60.97 | 61.82 | 57.85 | |
| LLaMA-2-13B | VersaTune Constant | 46.26↓ | 41.72↓ | 38.90↓ | 56.12↓ | 42.75↓ | 56.09↓ | 46.97↓ |
| VersaTune | 48.01 | 46.73 | 47.30 | 63.36 | 46.15 | 57.14 | 51.45 | |
| Qwen-2.5-14B | VersaTune Constant | 51.29↓ | 44.88↓ | 34.14↓ | 63.39↓ | 78.53↓ | 65.75↓ | 56.33↓ |
| VersaTune | 53.59 | 47.70 | 72.26 | 72.88 | 82.06 | 65.75 | 65.71 |
| Model | Method | LegalBench | LawBench |
| LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 |
| Inverse Distribution | 13.23↓ | 26.94↓ | |
| VersaTune Constant | 21.47↑ | 35.55↑ | |
| VersaTune | 23.18↑ | 36.31↑ | |
| LLaMA-3-8B | Uniform Distribution | 33.52 | 31.16 |
| Inverse Distribution | 27.83↓ | 27.48↓ | |
| VersaTune Constant | 47.85↑ | 37.75↑ | |
| VersaTune | 49.67↑ | 37.87↑ | |
| Qwen-1.5-7B | Uniform Distribution | 30.33 | 33.49 |
| Inverse Distribution | 29.80↓ | 30.13↓ | |
| VersaTune Constant | 37.97↑ | 35.52↑ | |
| VersaTune | 40.04↑ | 35.95↑ | |
| Qwen-2-7B | Uniform Distribution | 39.05 | 31.99 |
| Inverse Distribution | 34.01↓ | 27.81↓ | |
| VersaTune Constant | 45.86↑ | 32.72↑ | |
| VersaTune | 50.56↑ | 35.54↑ | |
| Qwen-2.5-7B | Uniform Distribution | 40.11 | 31.48 |
| Inverse Distribution | 36.36↓ | 26.98↓ | |
| VersaTune Constant | 48.78↑ | 35.20↑ | |
| VersaTune | 51.65↑ | 36.75↑ | |
| LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 |
| Inverse Distribution | 40.12↓ | 30.67↓ | |
| VersaTune Constant | 53.79↑ | 38.73↑ | |
| VersaTune | 55.87↑ | 40.14↑ | |
| Qwen-2.5-14B | Uniform Distribution | 50.73 | 39.49 |
| Inverse Distribution | 46.08↓ | 35.36↓ | |
| VersaTune Constant | 56.94↑ | 45.64↑ | |
| VersaTune | 60.59↑ | 46.58↑ |
| Model | Method | MedQA | MedMCQA |
| LLaMA-2-7B | Uniform Distribution | 23.45 | 27.57 |
| Inverse Distribution | 21.38↓ | 26.52↓ | |
| VersaTune Constant | 30.17↑ | 36.72↑ | |
| VersaTune | 35.04↑ | 40.75↑ | |
| LLaMA-3-8B | Uniform Distribution | 31.03 | 10.26 |
| Inverse Distribution | 25.51↓ | 8.77↓ | |
| VersaTune Constant | 37.33↑ | 30.15↑ | |
| VersaTune | 42.21↑ | 45.72↑ | |
| Qwen-1.5-7B | Uniform Distribution | 21.27 | 12.14 |
| Inverse Distribution | 16.15↓ | 9.91↓ | |
| VersaTune Constant | 25.60↑ | 30.28↑ | |
| VersaTune | 30.47↑ | 32.24↑ | |
| Qwen-2-7B | Uniform Distribution | 35.07 | 17.73 |
| Inverse Distribution | 23.90↓ | 16.31↓ | |
| VersaTune Constant | 40.89↑ | 39.13↑ | |
| VersaTune | 45.48↑ | 41.24↑ | |
| Qwen-2.5-7B | Uniform Distribution | 25.17 | 25.84 |
| Inverse Distribution | 24.16↓ | 19.35↓ | |
| VersaTune Constant | 30.20↑ | 49.71↑ | |
| VersaTune | 34.28↑ | 52.09↑ | |
| LLaMA-2-13B | Uniform Distribution | 32.98 | 36.54 |
| Inverse Distribution | 26.27↓ | 28.78↓ | |
| VersaTune Constant | 40.69↑ | 42.74↑ | |
| VersaTune | 45.78↑ | 47.67↑ | |
| Qwen-2.5-14B | Uniform Distribution | 47.85 | 38.71 |
| Inverse Distribution | 45.75↓ | 32.56↓ | |
| VersaTune Constant | 48.11↑ | 41.64↑ | |
| VersaTune | 50.24↑ | 45.15↑ |
| Model | Method | FinEval | FinanceIQ |
| LLaMA-2-7B | Uniform Distribution | 33.50 | 2.71 |
| Inverse Distribution | 32.96↓ | 2.53↓ | |
| VersaTune Constant | 35.89↑ | 6.28↑ | |
| VersaTune | 36.27↑ | 29.04↑ | |
| LLaMA-3-8B | Uniform Distribution | 34.83 | 4.97 |
| Inverse Distribution | 33.71↓ | 3.31↓ | |
| VersaTune Constant | 37.93↑ | 25.27↑ | |
| VersaTune | 38.80↑ | 43.58↑ | |
| Qwen-1.5-7B | Uniform Distribution | 53.87 | 14.17 |
| Inverse Distribution | 53.80↓ | 12.44↓ | |
| VersaTune Constant | 55.06↑ | 37.52↑ | |
| VersaTune | 55.52↑ | 56.63↑ | |
| Qwen-2-7B | Uniform Distribution | 59.49 | 14.62 |
| Inverse Distribution | 56.53↓ | 11.30↓ | |
| VersaTune Constant | 60.63↑ | 40.82↑ | |
| VersaTune | 60.95↑ | 68.39↑ | |
| Qwen-2.5-7B | Uniform Distribution | 59.58 | 31.66 |
| Inverse Distribution | 57.07↓ | 29.25↓ | |
| VersaTune Constant | 62.94↑ | 48.47↑ | |
| VersaTune | 62.48↑ | 69.09↑ | |
| LLaMA-2-13B | Uniform Distribution | 37.54 | 32.85 |
| Inverse Distribution | 36.67↓ | 26.76↓ | |
| VersaTune Constant | 39.33↑ | 38.47↑ | |
| VersaTune | 39.48↑ | 55.12↑ | |
| Qwen-2.5-14B | Uniform Distribution | 64.72 | 64.39 |
| Inverse Distribution | 64.88↑ | 60.53↓ | |
| VersaTune Constant | 65.03↑ | 73.24↑ | |
| VersaTune | 65.84↑ | 78.68↑ |
| Model | Method | SciEval | MMLU-Sci |
| LLaMA-2-7B | Uniform Distribution | 9.30 | 42.89 |
| Inverse Distribution | 8.98↓ | 39.67↓ | |
| VersaTune Constant | 49.91↑ | 45.87↑ | |
| VersaTune | 56.75↑ | 50.06↑ | |
| LLaMA-3-8B | Uniform Distribution | 6.51 | 50.17 |
| Inverse Distribution | 6.09↓ | 46.62↓ | |
| VersaTune Constant | 54.77↑ | 56.04↑ | |
| VersaTune | 56.67↑ | 60.61↑ | |
| Qwen-1.5-7B | Uniform Distribution | 7.63 | 50.67 |
| Inverse Distribution | 6.54↓ | 48.52↓ | |
| VersaTune Constant | 38.39↑ | 57.33↑ | |
| VersaTune | 47.76↑ | 61.11↑ | |
| Qwen-2-7B | Uniform Distribution | 25.30 | 62.73 |
| Inverse Distribution | 18.57↓ | 58.25↓ | |
| VersaTune Constant | 45.93↑ | 67.29↑ | |
| VersaTune | 51.58↑ | 70.42↑ | |
| Qwen-2.5-7B | Uniform Distribution | 19.88 | 65.84 |
| Inverse Distribution | 16.68↓ | 62.78↓ | |
| VersaTune Constant | 56.04↑ | 71.96↑ | |
| VersaTune | 68.14↑ | 74.16↑ | |
| LLaMA-2-13B | Uniform Distribution | 45.72 | 50.77 |
| Inverse Distribution | 38.96↓ | 48.68↓ | |
| VersaTune Constant | 57.13↑ | 55.10↑ | |
| VersaTune | 63.87↑ | 62.84↑ | |
| Qwen-2.5-14B | Uniform Distribution | 39.74 | 73.45 |
| Inverse Distribution | 27.68↓ | 68.22↓ | |
| VersaTune Constant | 48.31↑ | 78.46↑ | |
| VersaTune | 62.89↑ | 82.86↑ |
| Model | Method | HumanEval | MBPP |
| LLaMA-2-7B | Uniform Distribution | 5.67 | 3.44 |
| Inverse Distribution | 3.47↓ | 2.42↓ | |
| VersaTune Constant | 12.47↑ | 14.47↑ | |
| VersaTune | 15.62↑ | 15.68↑ | |
| LLaMA-3-8B | Uniform Distribution | 22.94 | 28.85 |
| Inverse Distribution | 19.67↓ | 24.34↓ | |
| VersaTune Constant | 29.88↑ | 33.22↑ | |
| VersaTune | 28.91↑ | 35.65↑ | |
| Qwen-1.5-7B | Uniform Distribution | 32.86 | 25.86 |
| Inverse Distribution | 28.55↓ | 23.94↓ | |
| VersaTune Constant | 38.41↑ | 29.67↑ | |
| VersaTune | 40.49↑ | 31.80↑ | |
| Qwen-2-7B | Uniform Distribution | 53.26 | 37.82 |
| Inverse Distribution | 50.65↓ | 33.63↓ | |
| VersaTune Constant | 56.71↑ | 45.87↑ | |
| VersaTune | 58.15↑ | 47.64↑ | |
| Qwen-2.5-7B | Uniform Distribution | 55.64 | 46.86 |
| Inverse Distribution | 52.97↓ | 44.63↓ | |
| VersaTune Constant | 59.15↑ | 52.10↑ | |
| VersaTune | 60.68↑ | 61.25↑ | |
| LLaMA-2-13B | Uniform Distribution | 36.54 | 38.55 |
| Inverse Distribution | 28.78↓ | 35.83↓ | |
| VersaTune Constant | 42.74↑ | 42.76↑ | |
| VersaTune | 47.67↑ | 44.62↑ | |
| Qwen-2.5-14B | Uniform Distribution | 68.75 | 72.14 |
| Inverse Distribution | 63.36↓ | 68.49↓ | |
| VersaTune Constant | 78.72↑ | 78.33↑ | |
| VersaTune | 82.64↑ | 81.48↑ |
| Model | Method | AGIEval | HellaSwag |
| LLaMA-2-7B | Uniform Distribution | 20.16 | 71.40 |
| Inverse Distribution | 18.83↓ | 71.33↓ | |
| VersaTune Constant | 22.31↑ | 71.89↑ | |
| VersaTune | 24.67↑ | 71.76↑ | |
| LLaMA-3-8B | Uniform Distribution | 23.87 | 73.26 |
| Inverse Distribution | 23.45↓ | 72.40↓ | |
| VersaTune Constant | 25.62↑ | 73.33↑ | |
| VersaTune | 28.78↑ | 73.62↑ | |
| Qwen-1.5-7B | Uniform Distribution | 39.19 | 69.80 |
| Inverse Distribution | 39.56↑ | 69.59↓ | |
| VersaTune Constant | 42.50↑ | 70.67↑ | |
| VersaTune | 42.96↑ | 71.51↑ | |
| Qwen-2-7B | Uniform Distribution | 47.31 | 73.60 |
| Inverse Distribution | 45.74↓ | 73.52↓ | |
| VersaTune Constant | 48.16↑ | 72.98↓ | |
| VersaTune | 48.02↑ | 73.67↑ | |
| Qwen-2.5-7B | Uniform Distribution | 45.42 | 73.69 |
| Inverse Distribution | 45.67↑ | 72.92↓ | |
| VersaTune Constant | 47.75↑ | 73.88↑ | |
| VersaTune | 49.73↑ | 73.90↑ | |
| LLaMA-2-13B | Uniform Distribution | 36.89 | 73.50 |
| Inverse Distribution | 36.67↓ | 73.11↓ | |
| VersaTune Constant | 37.91↑ | 74.27↑ | |
| VersaTune | 39.64↑ | 74.63↑ | |
| Qwen-2.5-14B | Uniform Distribution | 54.92 | 75.88 |
| Inverse Distribution | 54.87↓ | 75.42↓ | |
| VersaTune Constant | 55.04↑ | 76.45↑ | |
| VersaTune | 55.52↑ | 75.98↑ |
D.2 Detailed Results in Domain Expansion Scenario
In this section, we exhibit the performance of the target domain and other non-target domains under the domain expansion scenario, as realized by Algorithm 3. The stacked group bar chart (left) illustrates the percentage change in performance for non-target domains relative to their pre-fine-tuning states, with the positive direction on the y-axis indicating performance improvement and the negative direction signifying a decline. The line chart (right) represents the overall change across all non-target domains for three distinct strategies, with color legends corresponding to those of the line chart on the right. The right-side chart depicts the percentage increase in performance for the current target domain. We employed the LLaMA-3-8B and LLaMA-2-13B models, and the mean percentage change in model performance when focusing on domain enhancement is presented in both the stacked group bar chart and the line chart.