### v4

# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

Abstract.
Large-scale pretrained models, particularly Large Language Models (LLMs), have exhibited remarkable capabilities in handling multiple tasks across domains due to their emergent properties. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-ability performances during training. We categorize knowledge into distinct domains including law, medicine, finance, science, code, etc. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the training data composition that aligns with the model’s existing knowledge distribution. During the training process, domain weights are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results demonstrate that VersaTune achieves significant improvements in multi-domain performance, with an enhancement in comprehensive multi-domain tasks. Additionally, in scenarios where specific domain optimization is required, VersaTune reduces the degradation of performance in other domains by , while preserving the target domain’s training efficacy.
PVLDB Reference Format:
Keer Lu, Keshi Zhao, Zheng Liang, Da Pan, Shusen Zhang, Xin Wu, Zenan Zhou, Guosheng Dong, Bin Cui, Wentao Zhang. PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX
† Corresponding Author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and other artifacts will be released after the paper is accepted.
1. Introduction
In recent years, large-scale pretrained models with enormous pretraining tokens and parameters, such as Transformer models, have become a cornerstone in Artificial Intelligence (AI) (Achiam et al., 2023; Dwivedi et al., 2021; Lewkowycz et al., 2022), particularly for Natural Language Processing tasks (Brown et al., 2020; Devlin, 2018; Radford et al., 2019), reshaping AI research and applications in domains such as law (Cui et al., 2023), medicine (Singhal et al., 2023; Thirunavukarasu et al., 2023), finance (Li et al., 2023b; Wu et al., 2023), science (Beltagy et al., 2019; Taylor et al., 2022) and code (Liu et al., 2024; Roziere et al., 2023). The success of Transformer models, such Large Language Models (LLMs), stems from their capabilities to automatically learn and distill hierarchical data representations, making them highly effective for complex tasks (Nie et al., 2023). In order to further enhance such abilities across these areas, LLMs typically undergo the supervised fine-tuning (SFT) stages on domain-specific datasets.
Data plays a fundamental role throughout the training phases of LLMs. While the significance of model architectures is widely acknowledged, data-centric approaches hold substantial importance (Wang et al., 2023), which emphasizes the quality, organization, and management of data as the crucial factor for improving model performance (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023). Efficient data management, especially in constructing an appropriately tailored training dataset with suitable data composition, is critical for synergizing the model’s performance and ability to generalize across different tasks and domains during both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017).
As demonstrated by the robust performances of state-of-the-art LLMs such as GPT-4 (Achiam et al., 2023), PaLM (Anil et al., 2023) and Gemini (Team et al., 2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on supervised fine-tuning tends to merely concentrate on a single ability of LLMs (Dong et al., 2023; Xu et al., 2024), with the multi-task performance on composite data of essentially different downstream tasks being less studied. We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during training:
How to design a data composition strategy that could achieve overall multitasking capabilities?
Through extensive experimental analysis, we identified that the challenges associated with data composition strategies stem from the following three key aspects:
C1: Catastrophic Forgetting. Given the fundamental differences between tasks across various domains, for multi-domain SFT, the sequential training strategy of specific datasets in multiple phases can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting (Kaushik et al., 2021; McCloskey and Cohen, 1989), as illustrated in Table 1 and Figure 2. This problem hinders the versatile fine-tuning performance of a model across multiple domains (De Lange et al., 2021; Dong et al., 2023; Yuan et al., 2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during the SFT stage, enhancing the overall performance and adaptability.
C2: Low Efficiency. Existing research on data composition during the supervised fine-tuning phase for LLMs is still in its initial stages, with most strategies based on heuristic or manually determined rules (Albalak et al., 2024; Dubey et al., 2024; Wang et al., 2023). One of the common baselines is defining domain weights referring to the natural domain sizes, which weights all the individual data points equally. Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience (Azerbayev et al., 2023; Roziere et al., 2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of different domains during the training process, which does not allow for real-time feedback from the LLMs in the current phase to inform and optimize the data composition. This static approach hinders the minimization of performance loss in other domains as LLMs undergo specialized training.
To address these challenges, we introduce VersaTune, a novel data composition framework designed for enhancing models’ overall performances across different vertical domains during supervised fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 3.1), followed by data composition based on the existing distribution for multi-ability enhancement (Section 3.2.2) as well as flexible domain expansion (Section 3.2.3). VersaTune exhibits the following properties:
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ fine-tuning stage, providing an efficient data composition strategy for enhancing versatile capabilities (for C2).
Flexible. VersaTune can be flexibly adapted to scenarios that expand performance on specific domain tasks while minimizing the degradation of the model’s capabilities in non-target domains (for C1, C3).
Robust. The data composition strategy demonstrates significant improvements after fine-tuning in open-source models with parameter sizes ranging from 7B-14B, adding to the effectiveness of VersaTune (for C1, C2 and C3).
In summary, our contributions are as follows:
Challenges in Multi-Capability Training. Through a quantitative analysis of the performance losses in non-target domains resulting from existing training methods, we emphasize the challenge of data composition in multi-domain task training (Section 2).
Multi-Capability Data Composition Framework. We introduce VersaTune, a data composition framework that leverages the intrinsic domain knowledge distribution of the model to proportion training data. VersaTune is designed to enhance overall performance across multiple domains while also providing flexibility for targeted improvements in specific domains (Section 3).
Performance and Effectiveness. Our comprehensive evaluations across domains demonstrate that VersaTune can achieve an improvement of in versatile fine-tuning for multiple domains. Furthermore, when focusing on specific-domain fine-tuning, VersaTune maintains the training effectiveness in the target domain while reducing the performance degradation in other non-target domains by (Section 5).
2. Preliminary
2.1. Transformer Models
Transformers were initially introduced to address sequence modeling and transduction challenges, including language modeling and machine translation (Vaswani, 2017). The core of each Transformer layer consists of self-attention mechanisms and point-wise feed-forward networks. Self-attention network enables the model to assess the importance of various words in a sequence regardless of their positions (Zhang et al., 2019). It calculates a score for each word, indicating the degree of attention to be directed towards other sequence parts during encoding. This process involves query, key, and value weight sets that are optimized through training. Feed-forward network (FFN) applies the same linear transformation with a ReLU activation to each position individually. The FFN’s role is to independently transform each position’s representation, introducing non-linearity into the model and allowing it to learn complex functions (Miao et al., 2022). This component complements the self-attention mechanism by processing each input position in isolation, thus enhancing the model’s adaptability to diverse data types and tasks. These operations, predominantly dense algebraic computations such as matrix multiplications, lead to significant computational expenses and memory demands.
Domain of Training Dataset | Training Step | Variations in Comprehensive Domains () | Sum. () | |||||
(Epoch) | Law | Medicine | Finance | Science | Code | Other | ||
Law | 1 | - | 18.82 | 14.71 | 11.76 | 11.18 | 5.00 | 32.05 |
2 | - | 12.65 | 30.59 | 4.41 | 5.29 | 11.76 | 3.52 | |
3 | - | 17.94 | 12.35 | 8.82 | 23.53 | 5.00 | 42.94 | |
4 | - | 5.00 | 2.06 | 31.18 | 21.76 | 12.65 | 72.65 | |
Medicine | 1 | 10.29 | - | 3.82 | 24.12 | 19.41 | 7.35 | 16.75 |
2 | 18.82 | - | 6.47 | 40.00 | 7.94 | 17.65 | 10.88 | |
3 | 22.35 | - | 8.82 | 7.94 | 19.12 | 10.00 | 52.35 | |
4 | 27.94 | - | 11.76 | 2.35 | 21.76 | 12.65 | 76.46 | |
Finance | 1 | 20.59 | 7.94 | - | 10.29 | 12.65 | 6.47 | 16.76 |
2 | 18.24 | 9.71 | - | 9.41 | 5.29 | 8.82 | 4.41 | |
3 | 23.53 | 9.41 | - | 17.35 | 14.71 | 7.94 | 25.88 | |
4 | 5.00 | 9.12 | - | 20.29 | 12.94 | 21.76 | 59.11 | |
Science | 1 | 10.29 | 17.06 | 3.82 | - | 4.71 | 7.35 | 9.11 |
2 | 11.47 | 12.35 | 4.71 | - | 5.88 | 12.94 | 22.65 | |
3 | 21.76 | 7.94 | 8.82 | - | 4.41 | 10.00 | 37.05 | |
4 | 27.94 | 2.35 | 11.47 | - | 12.65 | 12.59 | 62.30 | |
Code | 1 | 3.82 | 7.35 | 17.35 | 9.12 | - | 7.29 | 11.99 |
2 | 9.71 | 6.47 | 7.94 | 5.29 | - | 6.18 | 25.01 | |
3 | 22.35 | 8.82 | 14.12 | 7.94 | - | 10.02 | 47.37 | |
4 | 26.18 | 7.06 | 8.82 | 3.24 | - | 22.65 | 67.95 |
2.2. LLMs Training
Large Language Models (LLMs) are Transformer-based models containing an extensive parameter count in the billions or greater, which are trained on a vast scale of datasets (Zhao et al., 2023b). The training process of LLMs generally involves the pretraining and fine-tuning stages. We have outlined several concepts about LLMs training.
2.2.1. Pretraining
Large Language Models (LLMs) establish basic knowledge abilities, including language understanding and text generation, during the pretraining stage (Brown et al., 2020). In this stage, LLMs engage in unsupervised training through the processing of extensive raw text corpora, thereby enhancing their capabilities in language modeling. For a given sequence , the typical task for LLMs involves the prediction of the subsequent token given the preceding tokens as contextual input. The goal is to maximize the likelihood function presented in Equation 1:
Beginning of Sequence (¡BOS¿) During the above process, the Beginning of Sequence (¡BOS¿) token plays an important role, which serves as a signal to the model that the input sequence is starting (Brown et al., 2020; Li et al., 2023a; Wu et al., 2025). It can be thought of as a special marker that indicates the start of a new sequence, allowing the model to reset its context and begin processing a new piece of text. In the context of pretraining, the ¡BOS¿ token is used to initialize the input to the model, and it can be concatenated with the actual text data to form the input sequence. This token helps the model to differentiate between the start of a new input and the continuation of an existing one. It is particularly crucial in tasks where the model needs to generate text or understand the beginning of a new sentence or document, which helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on. The use of ¡BOS¿ tokens, along with other special tokens like ¡EOS¿ (End of Sequence), helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on.
2.2.2. Supervised Fine-Tuning
The Supervised Fine-Tuning (SFT) stage involves further training on the pretrained model to refine it task-solving capabilities and ensure greater alignment with human instructions across multiple domains (Zhao et al., 2023b; Nie et al., 2023). During the SFT phase, it is a typical practice to employ datasets specific to a particular domain for the fine-tuning of LLMs, which may lead to a decline in performance on non-target domains, a phenomenon commonly referred to as catastrophic forgetting. We conducted experiments on open-sourced models including LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series to assess how the model’s proficiency in other domains changes when fine-tuned with data from a single domain, as illustrated in Table 1 and Figure 2. We have regulated the number of training instances per epoch to a fixed count of 10,000. More details about training and evaluation settings can be found in Section 5.1. Our findings indicate that when a model is trained exclusively with data from a single domain, its performance on tasks from other domains tends to degrade progressively over the course of training.
While recent research has delved into exploring fine-tuning methods for multi-task enhancement (Dong et al., 2023; Sanh et al., 2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage to bridge this gap.
3. VersaTune
In this section, we introduce VersaTune, a novel data composition framework designed for multi-capability training. This framework aims to effectively compose data from multiple domains and optimize the data composition during training. Figure 3 presents the workflow of VersaTune, which generally contains two phases.
3.1. Phase 1: Knowledge Distribution Detection
Here, we first present a domain mixing strategy for fine-tuning a LLM that possesses a comprehensive multitask capability (Section 3.1.1). This approach is designed to align with the inherent domain knowledge distribution within the base model waiting for subsequent training. Following this, we describe the method for detecting domain knowledge proportion of the base model, which is crucial for informing the fine-tuning process (Section 3.1).
3.1.1. Knowledge Consistency Training
Previous research on data mixing ratios during the SFT phase for LLMs has predominantly focused on enhancing capabilities within a specific domain, often utilizing only data from that domain or employing heuristic, experience-based data proportions. We argue that such data mixing strategies can significantly impair the LLM’s abilities in other domains. In the fine-tuning stage, maintaining a robust overall capability across various domains is crucial.
What data mixing strategy effectively boosts the versatile performance of LLMs across different domains during the SFT phase? We propose the following statement:
Statement 1 (Knowledge Consistency Training). An LLM fine-tuned with domain-specific data proportions that align with its pretrained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
where denotes the set of all possible data points.
The rationale behind this statement is rooted in the observation that during the pretraining phase, LLMs develop a general understanding of language features and domain-specific knowledge. By maintaining the same distribution of knowledge during fine-tuning, the model can build upon this pre-existing knowledge, thereby enhancing learning efficiency and robustness.
3.1.2. Domain Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods (Gekhman et al., 2024; Zhao et al., 2023a) and training data inference strategies for LLMs (Ding et al., 2022; Hayase et al., 2024), we propose a structured approach to efficiently detect domain knowledge based on statistics. The process involves the generation of textual outputs from the base model waiting to be fine-tuning, followed by their classification into predefined domains referring to a proprietary model . The above process is repeated multiple times to ensure statistical robustness.
Assuming the data corpus encompasses distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary model to infer the probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for the domain knowledge distribution.
Input: Base model to be fine-tuned , Proprietary model , Hyperparameters: sample number , maximum iterations
Parameter: Data samples generated from the base model
Output: Domain distribution
Define : domain probability distribution of data sample
3.2. Phase 2: Fine-Tuning Multi-Ability LLMs Efficiently
Having detected the distribution of domain knowledge within the base model, we will now utilize these findings to guide our multi-ability SFT process. The approaches are designed to enhance the overall performance of the fine-tuned model across a spectrum of multi-domain tasks (Section 3.2.2), as well as to facilitate the flexible expansion of capabilities in specific domains (Section 3.2.3).
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a specific domain and its corresponding proportion, such that training a model on dataset could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performance of other domains. The specialized capabilities of LLMs are ultimately measured using downstream tasks related to different domains (e.g., FinBen (Xie et al., 2024a) for financial performances).
3.2.1. Learnable Potential and Forgetting Degree of Knowledge
Before formally introducing the effective multi-task fine-tuning and flexible domain expansion data composing strategies, we will first provide an overview of the evaluation metrics used for the following algorithms in this subsection.
Mastery Ceiling. We first fine-tuned the small reference model for epochs on each domain separately, and identified the epoch with the lowest average loss during this process as the lower bound on the minimum loss attainable by the target model for the given domain. This value represents the highest level of domain knowledge mastery that the model can achieve in the context of the current specific domain under given conditions.
Learnable Potential. We can observe whether a domain could be effectively learned by the model through comparing the difference between the loss of the target model and the minimum loss that the reference model can achieve. Based on these principles, we propose Equation 3 to score the learnable potential of domain :
where denotes the loss associated with the target model for the -th domain, while signifies the corresponding loss for the reference model within the same domain. To mitigate the impact of inherent loss variations across different domains for the model, we have introduced a normalization term into the formula.
Forgetting Degree. When focusing on expanding a model to a specific domain, our objective is to minimize the loss of the model’s knowledge regarding other domains. Here we segment the fine-tuning stage into distinct checkpoints. We quantify the degree of knowledge loss, or the forgetting of the current domain, by measuring the difference in loss between the -th and -th training steps. This difference reflects the model’s mastery loss for the tasks associated with the current domain. Based on this principle, we introduce Equation 4 to assess the model’s forgetting degree for domain at the -th training step.
where represents the loss at the -th training step associated with the target model for the -th domain, while denotes the loss at the preceding -th iteration for the same domain. We also incorporated a normalization factor into the equation to counteract the effects of inherent loss disparities among domains.
3.2.2. Effective Multi-Ability Fostering
Input: Base model to be fine-tuned , Domain reference loss , Hyperparameters: magnitude of adjustment , number of training steps
Parameter: Data proportion of the SFT dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
To cultivate the multi-tasking capabilities of a LLM during the fine-tuning phase, we have aligned the initial domain distribution in the SFT stage with the knowledge detection results of the base model as stated in Equation 2. Furthermore, we dynamically make minor adjustments in the composition ratios of various domains based on the model’s real-time feedback at different SFT stages.
As detailed in Algorithm 2, in the pursuit of balanced domain expertise enhancement, we first blended the domain proportions in accordance with the base model’s intrinsic domain knowledge distribution detected by Algorithm 1 (Line 1). Then at each training step , we assigned a learnable potential score to each domain based on the methodology outlined in Equation 3. These scores were then utilized to fine-tune the representation of each domain within the composite SFT dataset, ensuring a balanced development of competencies across all domains throughout the fine-tuning process (Line 3-7). At the same time, the parameters of model are updated based on the gradients computed through backpropagation (Line 8). This adaptive approach is imperative to harmonize the progression of capabilities in different domains and to optimize the model’s performance on a multifaceted array of tasks.
3.2.3. Flexible Domain Expansion
Input: Base model to be fine-tuned , Domains that require enhanced cultivation , Domain reference loss , Hyperparameters: number of training steps , magnitude of adjustment , extent of domain proportion adjustment , variation threshold
Parameter: Data proportion of the SFT dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
Define : forgetting degree of the current domain
When conducting fine-tuning on a pretrained model, there are instances where we aim to particularly enhance the model’s performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
Statement 2
When fine-tuning a LLM for a specific capability, increasing the volume of data from a particular domain while adjusting other domains according to the knowledge distribution of the base model, facilitates a flexible strategy for domain expansion. Formally, the relationship can be represented as:
where is the data distribution in the given domain , and is the adjustment factor.
Here is determined based on the knowledge distribution of the pre-trained domain. In particular, when increases, the other domains shrink proportionally as a whole, which can be expressed as:
where is the increased adjustment factor, and is the original ratio of other domain knowledge relative to .
As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pretrained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current training phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other non-target domains: if the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other non-target domains according to Equation 6 (Line 8-9); otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
4. Implementation
VersaTune offers a comprehensive solution for data composition aimed at enhancing performance across multiple domains. In this section, we will briefly discuss the implementation of key components and the setting of hyperparameters.
Domain Knowledge Detection. During the knowledge distribution detection phase for our target models, we employed Qwen2.5-72B-Instruct111https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as the proprietary model . For each target model slated for fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 4. The pre-existing domain knowledge distribution varies among different models. Therefore, it is essential to develop a data composition strategy that is tailored to the specific model being trained.
Hyper-Parameters Setting. During the multi-domain task fine-tuning of LLMs, we configured the number of training steps in Algorithm 2 and Algorithm 3 to 4 epochs. We experimented with various magnitude of adjustment, specifically , and observed consistent weight ordering across domains, which far outperformed our baselines (detailed in Section 5.2). Based on these experimental outcomes, we set the magnitude of adjustment to . Additionally, in the context of domain expansion, we set the increment for the target domain to per training step, considering the overall domain weight distribution across models. The variation threshold, denoted as , reflects the trade-off between enhancing specific domain skills and mitigating the loss of capabilities in non-target domains, where we assigned a weight of .
5. Experiments
Dataset | # Instance | Source | # Rounds | Full |
Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset222https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
MedQA | 10178 | Professional Medical Board Exams | 1 | Training Portion |
Finance Alpaca | 68912 | Alpaca, FiQA, 1.3k Pairs Generated using GPT3.5 | 1 | |
Sonnet3.5 Science Conversations | 8835 | Scientific Conversations with Sonnet3.5 | 11.1 | |
Code Alpaca | 20022 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | |
Alpaca | 49,087 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | Excluding Samples of Other Domains |
Domain | Benchmark | # Instance | Language | N-Shot |
Law | LegalBench (Guha et al., 2024) | 90,394 (164 sub-tasks) | English | 1 |
LawBench (Fei et al., 2023) | 10,000 (20 sub-tasks) | Chinese | 1 | |
Medicine | MedQA (Jin et al., 2020) | 1,273 | English | 1 |
MedMCQA (Pal et al., 2022) | 4,183 | English | 1 | |
Finance | FinEval (Zhang et al., 2023) | 4,661 (34 sub-tasks) | Chinese | 1 |
FinanceIQ (Zhang and Yang, 2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |
Science | SciEval (Sun et al., 2024) | 15,901 | English | 1 |
MMLU-Sci (Hendrycks et al., 2020) | 2,999 (14 sub-tasks) | English | 0 | |
Code | HumanEval (Chen et al., 2021) | 164 | English | 0 |
MBPP (Austin et al., 2021) | 974 | English | 0 | |
Other (General) | AGIEval (Zhong et al., 2023) | 8,062 (20 sub-tasks) | English, Chinese | 0 |
HellaSwag (Zellers et al., 2019) | 10,003 | English | 0 |
In this section, we first describe the details of our experimental setup (Section 5.1), including training datasets, models used for experimentation and benchmarks. We then outline the baseline methods we use for comparison (Section 5.2) and our experimental results to prove the effectiveness of VersaTune (Section 5.3).
Model | Method | Law | Medical | Finance | Science | Code | General | ||||||
LegalBench | LawBench | MedQA | MedMCQA | FinEval | FinanceIQ | SciEval | MMLU-Sci | HumanEval | MBPP | AGIEval | HellaSwag | ||
Baichuan-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
Inverse Distribution | 13.23↓ | 26.94↓ | 21.38↓ | 26.52↓ | 32.96↓ | 2.53↓ | 8.98↓ | 39.67↓ | 3.47↓ | 2.42↓ | 18.83↓ | 71.33↓ | |
VersaTune Constant | 21.47↑ | 35.55↑ | 30.17↑ | 36.72↑ | 35.89↑ | 6.28↑ | 49.91↑ | 45.87↑ | 12.47↑ | 14.47↑ | 22.31↑ | 71.89↑ | |
VersaTune | 23.18↑ | 36.31↑ | 35.04↑ | 40.75↑ | 36.27↑ | 29.04↑ | 56.75↑ | 50.06↑ | 15.62↑ | 15.68↑ | 24.67↑ | 71.76↑ | |
LLaMA-3-8B | Uniform Distribution | 33.52 | 31.16 | 31.03 | 10.26 | 34.83 | 4.97 | 6.51 | 50.17 | 22.94 | 28.85 | 23.87 | 73.26 |
Inverse Distribution | 27.83↓ | 27.48↓ | 25.51↓ | 8.77↓ | 33.71↓ | 3.31↓ | 6.09↓ | 46.62↓ | 19.67↓ | 24.34↓ | 23.45↓ | 72.40↓ | |
VersaTune Constant | 47.85↑ | 37.75↑ | 37.33↑ | 30.15↑ | 37.93↑ | 25.27↑ | 54.77↑ | 56.04↑ | 29.88↑ | 33.22↑ | 25.62↑ | 73.33↑ | |
VersaTune | 49.67↑ | 37.87↑ | 42.21↑ | 45.72↑ | 38.80↑ | 43.58↑ | 56.67↑ | 60.61↑ | 28.91↑ | 35.65↑ | 28.78↑ | 73.62↑ | |
Qwen-2-7B | Uniform Distribution | 39.05 | 31.99 | 35.07 | 17.73 | 59.49 | 14.62 | 25.30 | 62.73 | 53.26 | 37.82 | 47.31 | 73.60 |
Inverse Distribution | 34.01↓ | 27.81↓ | 23.90↓ | 16.31↓ | 56.53↓ | 11.30↓ | 18.57↓ | 58.25↓ | 50.65↓ | 33.63↓ | 45.74↓ | 73.52↓ | |
VersaTune Constant | 45.86↑ | 32.72↑ | 40.89↑ | 39.13↑ | 60.63↑ | 40.82↑ | 45.93↑ | 67.29↑ | 56.71↑ | 45.87↑ | 48.16↑ | 72.98↓ | |
VersaTune | 50.56↑ | 35.54↑ | 45.48↑ | 41.24↑ | 60.95↑ | 68.39↑ | 51.58↑ | 70.42↑ | 58.15↑ | 47.64↑ | 48.02↑ | 73.67↑ | |
Qwen-2.5-7B | Uniform Distribution | 40.11 | 31.48 | 25.17 | 25.84 | 59.58 | 31.66 | 19.88 | 65.84 | 55.64 | 46.86 | 45.42 | 73.69 |
Inverse Distribution | 36.36↓ | 26.98↓ | 24.16↓ | 19.35↓ | 57.07↓ | 29.25↓ | 16.68↓ | 62.78↓ | 52.97↓ | 44.63↓ | 45.67↑ | 72.92↓ | |
VersaTune Constant | 48.78↑ | 35.20↑ | 30.20↑ | 49.71↑ | 62.94↑ | 48.47↑ | 56.04↑ | 71.96↑ | 59.15↑ | 52.10↑ | 47.75↑ | 73.88↑ | |
VersaTune | 51.65↑ | 36.75↑ | 34.28↑ | 52.09↑ | 62.48↑ | 69.09↑ | 68.14↑ | 74.16↑ | 60.68↑ | 61.25↑ | 49.73↑ | 73.90↑ | |
Baichuan-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
Inverse Distribution | 40.12↓ | 30.67↓ | 26.27↓ | 28.78↓ | 36.67↓ | 26.76↓ | 38.96↓ | 48.68↓ | 28.78↓ | 35.83↓ | 36.67↓ | 73.11↓ | |
VersaTune Constant | 53.79↑ | 38.73↑ | 40.69↑ | 42.74↑ | 39.33↑ | 38.47↑ | 57.13↑ | 55.10↑ | 42.74↑ | 42.76↑ | 37.91↑ | 74.27↑ | |
VersaTune | 55.87↑ | 40.14↑ | 45.78↑ | 47.67↑ | 39.48↑ | 55.12↑ | 63.87↑ | 62.84↑ | 47.67↑ | 44.62↑ | 39.64↑ | 74.63↑ | |
Qwen-2.5-14B | Uniform Distribution | 50.73 | 39.49 | 47.85 | 38.71 | 64.72 | 64.39 | 39.74 | 73.45 | 68.75 | 72.14 | 54.92 | 75.88 |
Inverse Distribution | 46.08↓ | 35.36↓ | 45.75↓ | 32.56↓ | 64.88↑ | 60.53↓ | 27.68↓ | 68.22↓ | 63.36↓ | 68.49↓ | 54.87↓ | 75.42↓ | |
VersaTune Constant | 56.94↑ | 45.64↑ | 48.11↑ | 41.64↑ | 65.03↑ | 73.24↑ | 48.31↑ | 78.46↑ | 78.72↑ | 78.33↑ | 55.04↑ | 76.45↑ | |
VersaTune | 60.59↑ | 46.58↑ | 50.24↑ | 45.15↑ | 65.84↑ | 78.68↑ | 62.89↑ | 82.86↑ | 82.64↑ | 81.48↑ | 55.52↑ | 75.98↑ |
5.1. Experimental Setup
Training Datasets. We have collected datasets spanning 6 domains for supervised fine-tuning (SFT), including Lawyer-Instruct333https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA (Jin et al., 2020), Finance Alpaca444https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations555https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca666https://github.com/sahil280114/codealpaca and Alpaca (Taori et al., 2023), denoted as , to represent SFT datasets with respect to law, medicine, finance, science, code as well as general capabilities. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other specific five domains, keeping only the general domain instances unrelated to them. Details of the training datasets can be found in Table 2.
Model Details. We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b), Qwen (Bai et al., 2023; Yang et al., 2024) and Baichuan (Yang et al., 2023) series as our pretrained language models . For scenarios aimed at fostering multi-ability, we trained and assessed models including Baichuan-2-7B, LLaMA-3-8B, Baichuan-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. For domain expansion, the training and evaluations were performed using the Baichuan-2-7B and Baichuan-2-13B models. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Training Details. All experiments were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch, with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800.
Benchmarks. We evaluate the performance of the models on downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 3. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset (Zhang et al., 2024) for testing, aiming to ensure a robust and thorough evaluation.
5.2. Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. (3) VersaTune constant is implemented with fixed domain weights derived from the knowledge distribution obtained from probing the target model prior to fine-tuning, where we ablate the components of dynamic adaptation in Algorithm 2 for an in-depth analysis. Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
5.3. Results
We conduct evaluations to validate the efficiency of VersaTune across different open-source models in scenarios that encompass both effective multi-ability fostering and flexible domain expansion. We summarize the observations below.
Target Domain | Training Step | Method | Variations in Comprehensive Domains () | Sum. () | ||||||
(Epoch) | Law | Medicine | Finance | Science | Code | Other | Target | Non-Target | ||
Law | 100% Specific Domain | 5.89 | 18.82 | 14.71 | 11.76 | 11.18 | 5.00 | 5.89 | 32.05 | |
1 | Uniform Distribution of Non-Target Domains | 5.38 | 7.35 | 6.47 | 6.58 | 0.05 | 13.99 | 5.38 | 6.48 | |
VersaTune | 8.25 | 6.18 | 12.06 | 7.65 | 6.22 | 17.59 | 8.25 | 21.96 | ||
100% Specific Domain | 35.51 | 12.65 | 30.59 | 4.41 | 5.29 | 11.76 | 35.51 | 3.52 | ||
2 | Uniform Distribution of Non-Target Domains | 33.89 | 7.94 | 26.47 | 12.65 | 7.65 | 5.00 | 33.89 | 3.23 | |
VersaTune | 35.14 | 3.53 | 18.82 | 6.89 | 17.06 | 13.24 | 35.14 | 45.76 | ||
100% Specific Domain | 55.84 | 17.94 | 12.35 | 8.82 | 23.53 | 5.00 | 55.84 | 42.94 | ||
3 | Uniform Distribution of Non-Target Domains | 52.89 | 8.24 | 15.00 | 12.06 | 10.17 | 5.12 | 52.89 | 20.59 | |
VersaTune | 51.71 | 4.41 | 24.71 | 5.29 | 4.87 | 8.42 | 51.71 | 29.30 | ||
100% Specific Domain | 62.76 | 5.00 | 2.06 | 31.18 | 21.76 | 12.65 | 62.76 | 72.65 | ||
4 | Uniform Distribution of Non-Target Domains | 58.12 | 9.71 | 5.59 | 9.41 | 10.14 | 12.05 | 58.12 | 35.72 | |
VersaTune | 59.08 | 5.59 | 13.82 | 8.82 | 5.61 | 6.17 | 59.08 | 12.37 | ||
Medicine | 100% Specific Domain | 10.29 | 5.87 | 3.82 | 24.12 | 19.41 | 7.35 | 5.87 | 16.75 | |
1 | Uniform Distribution of Non-Target Domains | 3.82 | 5.36 | 5.29 | 5.59 | 2.65 | 6.53 | 5.36 | 0.36 | |
VersaTune | 3.53 | 8.17 | 7.65 | 8.82 | 16.47 | 6.17 | 8.17 | 15.00 | ||
100% Specific Domain | 18.82 | 40.44 | 6.47 | 40.00 | 7.94 | 17.65 | 40.44 | 10.88 | ||
2 | Uniform Distribution of Non-Target Domains | 7.94 | 33.68 | 9.12 | 20.59 | 4.98 | 5.64 | 33.68 | 2.87 | |
VersaTune | 12.35 | 35.21 | 7.65 | 8.84 | 16.52 | 12.65 | 35.21 | 58.01 | ||
100% Specific Domain | 22.35 | 58.78 | 8.82 | 7.94 | 19.12 | 10.00 | 58.78 | 52.35 | ||
3 | Uniform Distribution of Non-Target Domains | 10.46 | 55.69 | 14.98 | 12.35 | 4.41 | 14.18 | 55.69 | 31.68 | |
VersaTune | 3.53 | 53.85 | 4.52 | 17.06 | 7.64 | 6.03 | 53.85 | 22.68 | ||
100% Specific Domain | 27.94 | 64.61 | 11.76 | 2.35 | 21.76 | 12.65 | 64.61 | 76.46 | ||
4 | Uniform Distribution of Non-Target Domains | 13.82 | 58.07 | 11.18 | 2.35 | 6.47 | 18.53 | 58.07 | 47.65 | |
VersaTune | 5.59 | 59.81 | 4.27 | 10.68 | 5.94 | 10.26 | 59.81 | 15.38 | ||
Finance | 100% Specific Domain | 20.59 | 7.94 | 5.45 | 10.29 | 12.65 | 6.47 | 5.45 | 16.76 | |
1 | Uniform Distribution of Non-Target Domains | 12.05 | 7.36 | 8.21 | 6.74 | 19.41 | 8.53 | 8.21 | 8.83 | |
VersaTune | 15.07 | 4.12 | 10.46 | 3.24 | 17.06 | 8.55 | 10.46 | 22.70 | ||
100% Specific Domain | 18.24 | 9.71 | 34.97 | 9.41 | 5.29 | 8.82 | 34.97 | 4.41 | ||
2 | Uniform Distribution of Non-Target Domains | 20.59 | 7.94 | 31.08 | 4.71 | 7.69 | 4.98 | 31.08 | 5.23 | |
VersaTune | 24.70 | 7.35 | 33.92 | 5.02 | 7.08 | 10.58 | 33.92 | 44.69 | ||
100% Specific Domain | 23.53 | 9.41 | 55.87 | 17.35 | 14.71 | 7.94 | 55.87 | 25.88 | ||
3 | Uniform Distribution of Non-Target Domains | 15.02 | 8.24 | 52.41 | 12.06 | 10.30 | 4.98 | 52.41 | 20.56 | |
VersaTune | 24.71 | 11.18 | 53.04 | 5.29 | 4.44 | 8.83 | 53.04 | 32.09 | ||
100% Specific Domain | 5.00 | 9.12 | 62.89 | 20.29 | 12.94 | 21.76 | 62.89 | 59.11 | ||
4 | Uniform Distribution of Non-Target Domains | 5.88 | 5.29 | 56.13 | 14.09 | 13.23 | 20.87 | 56.13 | 47.60 | |
VersaTune | 14.19 | 8.24 | 58.47 | 13.24 | 8.52 | 9.41 | 58.47 | 25.22 | ||
Science | 100% Specific Domain | 10.29 | 17.06 | 3.82 | 6.78 | 4.71 | 7.35 | 6.78 | 9.11 | |
1 | Uniform Distribution of Non-Target Domains | 6.76 | 17.64 | 9.18 | 5.37 | 7.05 | 4.73 | 5.37 | 13.48 | |
VersaTune | 3.53 | 17.35 | 7.64 | 8.35 | 16.67 | 6.98 | 8.35 | 29.83 | ||
100% Specific Domain | 11.47 | 12.35 | 4.71 | 40.84 | 5.88 | 12.94 | 40.84 | 22.65 | ||
2 | Uniform Distribution of Non-Target Domains | 8.24 | 20.59 | 5.68 | 32.78 | 9.12 | 5.85 | 32.78 | 21.30 | |
VersaTune | 12.36 | 16.17 | 9.43 | 36.97 | 16.89 | 12.65 | 36.97 | 42.78 | ||
100% Specific Domain | 21.76 | 7.94 | 8.82 | 63.20 | 4.41 | 10.00 | 63.20 | 37.05 | ||
3 | Uniform Distribution of Non-Target Domains | 9.98 | 12.06 | 15.01 | 55.78 | 4.11 | 14.13 | 55.78 | 31.17 | |
VersaTune | 11.47 | 11.18 | 6.76 | 55.40 | 16.49 | 6.81 | 55.40 | 16.25 | ||
100% Specific Domain | 27.94 | 2.35 | 11.47 | 66.15 | 12.65 | 12.59 | 66.15 | 62.30 | ||
4 | Uniform Distribution of Non-Target Domains | 13.82 | 13.53 | 11.18 | 58.46 | 6.57 | 6.47 | 58.46 | 24.51 | |
VersaTune | 10.12 | 10.00 | 4.21 | 61.30 | 5.30 | 6.74 | 61.30 | 16.37 | ||
Code | 100% Specific Domain | 3.82 | 7.35 | 17.35 | 9.12 | 10.46 | 7.29 | 10.46 | 11.99 | |
1 | Uniform Distribution of Non-Target Domains | 3.76 | 5.09 | 7.80 | 7.68 | 5.23 | 5.65 | 5.23 | 6.86 | |
VersaTune | 5.06 | 11.82 | 8.76 | 8.96 | 5.98 | 8.19 | 5.98 | 15.15 | ||
100% Specific Domain | 9.71 | 6.47 | 7.94 | 5.29 | 47.28 | 6.18 | 47.28 | 25.01 | ||
2 | Uniform Distribution of Non-Target Domains | 7.90 | 13.49 | 9.05 | 12.03 | 38.31 | 17.76 | 38.31 | 9.19 | |
VersaTune | 12.22 | 15.04 | 6.78 | 16.28 | 39.77 | 6.14 | 39.77 | 32.02 | ||
100% Specific Domain | 22.35 | 8.82 | 14.12 | 7.94 | 61.95 | 10.02 | 61.95 | 47.37 | ||
3 | Uniform Distribution of Non-Target Domains | 9.96 | 15.07 | 8.46 | 5.33 | 55.62 | 10.04 | 55.62 | 38.20 | |
VersaTune | 17.39 | 17.25 | 9.09 | 12.31 | 56.12 | 6.19 | 56.12 | 9.27 | ||
100% Specific Domain | 26.18 | 7.06 | 8.82 | 3.24 | 64.76 | 22.65 | 64.76 | 67.95 | ||
4 | Uniform Distribution of Non-Target Domains | 14.01 | 13.86 | 6.57 | 6.66 | 58.06 | 13.93 | 58.06 | 41.71 | |
VersaTune | 5.66 | 4.42 | 10.10 | 9.95 | 59.71 | 13.57 | 59.71 | 14.96 |
VersaTune is efficient across different models in both scenarios. Table 4 and Figure 5 show that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 5 and Figure 6, where we averaged the experimental results from Baichuan-2-7B and Baichuan-2-13B. Three interesting phenomena are observed from the outcomes:
Absolute Count vs. Proportion. Notably, by the second epoch of training, the degradation in performance across non-target domains tends to be mitigated to some extent, and in some cases, there is even a positive trend in capability enhancement. We attribute this phenomenon to the fact that the absolute number of samples for each domain, relative to the distribution of domain samples, has a predominant influence on model performance at this stage.
Domain Interactions. Domains are not entirely orthogonal to each other, and there is a degree of mutual reinforcement among them. For instance, enhancing capabilities in the medicine domain can, to a certain degree, boost performance in the science domain, as illustrated in Figure 6.
Domain Mastery Efficiency. From the slope of the target domain performance increase in Figure 6 (b), it is evident that the model’s efficiency in mastering knowledge of a specific domain diminishes over training. In other words, as training progresses, the model’s grasp of the current target domain approaches saturation, while its performance on non-target domains declines sharply. Consequently, greater emphasis should be placed on mitigating losses in non-target domains during this phase, aiming to strike a balance between domain expansion and the salvage of capabilities in non-target domains, which is also demonstrated in Figure 7.
Knowledge consistency training boosts performance. In Table 4 and Figure 5, we also present the experimental results of data composition strategies that allocate domain data in a manner inversely proportional to the pre-existing knowledge distribution detected within each domain. As expected, the inverse distribution strategy yielded even lower performance compared to the simplest approach of uniform distribution, which evenly distributes data across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training, which proves the efficacy of knowledge consistency training stated in Section 3.1.1.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights referring to the knowledge distribution obtained from probing the target model prior to supervised fine-tuning, namely VersaTune Constant, to ablate the component of dynamic adaptation in Algorithm 2. Table 4 and Figure 5 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights throughout the training process by continuously monitoring the learnable potential within each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes during training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial.
Establishing proportion thresholds for specific domains counts during domain expansion. We consider conducting a comparative analysis between the outcomes of VersaTune and those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3, to ablate the component of criteria for determining the upper limit of domain expansion in Algorithm 3. In Figure 7, we present the trends in the overall multi-domain performance of the models under specific domain expansion for each domain. It can be observed that, for the majority of domains, the gap in average multi-task performance between models trained with VersaTune and those without an upper limit on domain proportion becomes increasingly pronounced after the second or third epoch. We speculate that it is because by the later stages of supervised fine-tuning, models’ proficiency in the target domain approaches saturation. Further increasing the proportion of the current domain provides diminishing returns and can lead to a significant loss of performance in other domains. Notably, between the second and third epochs of fine-tuning, the model reaches a balance where the efficiency of improvement in the target domain is matched by the rate of performance degradation in non-target domains. The finding shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Concurrently, it ensures gains in the capacity for the current domain of interest.
6. Related Work
Data Reweighting for LLM Training. Data reweighting maintains full access to the entire dataset while adjusts the relative importance of each instance for various target tasks, which is essential for both pretraining and fine-tuning stages of LLMs (Wang et al., 2023). During the pretraining stage, DoReMi (Xie et al., 2024b) and DoGE (Fan et al., 2023) employ lightweight proxy models to estimate weights for different data sources, which are subsequently applied to the formal training of LLMs. Furthermore, Sheared LLaMA (Xia et al., 2023) implements an online variant of DoReMi, where the loss reference provided by the proxy model is replaced with the state of the pre-trained model from the previous training step. Additionally, ODM (Albalak et al., 2023) leverages the differences in loss or perplexity across domains within a multi-armed bandit framework to dynamically adjust the data distribution along training steps. As for the SFT phase, Dong et al. (Dong et al., 2023) focus on enhancing the model’s math reasoning, coding, and general human-aligning abilities through a dual-stage mixed fine-tuning strategy. However, the mixing ratios for different domains rely heavily on enumeration and empirical methods, and the covered domains are not holistic. This paper provides a comprehensive overview of the model’s capabilities across various domains during the SFT stage and proposes appropriate multi-ability fine-tuning methods.
Knowledge Detection in LLMs. Investigating the knowledge contained in current LLMs is essential for guiding their subsequent training (Chang et al., 2024). The knowledge encompasses multiple dimensions, such as different domain sources and task attributes. Existing work in LLM knowledge detection primarily focuses on prompting and calibration. Directly prompting the model to generate sequences and extracting confidence scores from the model (Gekhman et al., 2024; Kadavath et al., 2022; Kuhn et al., 2023; Manakul et al., 2023) is a common strategy. However, such approaches are highly dependent on prompt design and task selection, which can introduce bias into the assessment. Other studies have attempted to infer the training data mixtures used in previous training stages of LLMs (Antoniades et al., 2024; Hayase et al., 2024; Hu et al., 2022; Ye et al., 2022). The essence of these studies is to evaluate the current knowledge state of the models and provide targeted strategies for data organization and management in subsequent training phases.
7. Conclusion
Large-scale pretrained models, such as Large Language Models (LLMs), have demonstrated exceptional performance across multiple domains due to their emergent properties. Versatile capability across multiple domains is of paramount importance, which is typically cultivated during the Supervised Fine-Tuning (SFT) phase of LLMs. However, existing data composition methods for the SFT stage of LLMs often focus on single domain tasks, leading to significant degradation in performance across other domains. Therefore, there is an urgent need for an effective data composition strategy during the SFT phase. In this paper, we introduce VersaTune, a data composition framework designed to enhance the multi-domain capabilities of models during the supervised fine-tuning phase, which is based on the domain knowledge distribution of the target model. Experimental results from different models and downstream tasks across various domains demonstrate that VersaTune achieves excellent training outcomes in both scenarios of overall multi-domain enhancement and flexible expansion of specific domains.